### v1

# Stochastic Multi-round Submodular Optimization with Budget

Abstract
In this work we study the problem of Stochastic Budgeted Multi-round Submodular Maximization (SBMSm), in which we would like to maximize the sum over multiple rounds of the value of a monotone and submodular objective function, subject to the fact that the values of this function depend on the realization of stochastic events and the number of observations that we can make over all rounds is limited by a given budget. This problem extends, and generalizes to multiple round settings, well-studied problems such as (adaptive) influence maximization and stochastic probing.
We first show that whenever a certain single-round optimization problem can be optimally solved in polynomial time, then there is a polynomial time dynamic programming algorithm that returns the same solution as the optimal algorithm, that can adaptively choose both which observations to make and in which round to have them. Unfortunately, this dynamic programming approach cannot be extended to work when the single-round optimization problem cannot be efficiently solved (even if we allow it would be approximated within an arbitrary small constant). Anyway, in this case we are able to provide a simple greedy algorithm for the problem. It guarantees a $(1/2-\epsilon)$-approximation to the optimal value, even if it non-adaptively allocates the budget to rounds.
1 Introduction
Consider the well-known problem of influence maximization: here we are given a (social) network, with edge probabilities encoding the strength of the relationships among network’s components; these probabilities describe how easily influence (e.g., an information or an innovation) spreads over the network starting from a limited set of nodes, usually termed seeds; the goal is then to find the set of seeds that maximizes the number of “influenced” nodes. The non-adaptive version of the problem, in which all the seeds are chosen before that influence starts, has been widely studied, with numerous successful algorithms [26, 10, 7] and applications ranging from viral marketing [26], to election manipulation [33, 13, 8], to infective disease prevention [34]. Recently, focus moved to the adaptive variant of this problem (in which seeds can be added even when the influence is spreading from previously placed seeds) [22, 28, 14]. This variant happens to have strong relationship with some Stochastic Probing problems [3], in which there are $n$ items whose states are stochastically defined according to (not necessarily equal) independent distributions and the goal is to choose a feasible subset of items maximizing the value of a non-negative, monotone and submodular set function. Several stochastic probing settings can be also encountered in expert hiring [24], online dating and kidney exchange [23], Bayesian mechanism design [2], and sequentially posted pricing [9, 1].
Research both in influence maximization and stochastic probing focused almost exclusively on the single-round setting. Recently, there is an increasing interest in those cases in which the optimization is run in a multi-round setting. This is highlighted both by the increasing amount of literature on the topic (see Related Works), and by the numerous applications that are evidently deployed over multiple rounds, as it is the case, e.g., for many real-word marketing or electoral campaigns, or expert hiring sessions. Moreover, it is often the case that in these marketing, electoral, or hiring campaigns we are interested both in maximizing the total effectiveness of the campaign, and in limiting the total budget spent. Note that, in order to achieve this goal it may be insufficient to limit the budget spent at each round, since this may lead in general to very inefficient outcomes (see Remark 1 below for more details). However, this makes the problem much more complex: we do not only need to care about which action to take at each round, but also we need to care about how large must be the portion of the budget spent at that round.
In this work, we aim to extend the study of above cited problems (and their generalizations) by considering all these properties: multiple-rounds, total budget, and adaptivity.
Related Works.
Previous works indeed never consider all these three properties jointly.
The interest in multi-round variants of the above discussed problems is testified by the increasing amount of papers focusing on this aspect: for example, Sun et al. [29] considered an influence maximization problem over multiple and independent rounds of diffusion, where the goal is to maximize the number of nodes selected in at least one round, subject to the selection of at most $k$ seeds for each round (thus, in contrast to our setting, they have a limit only on per-round budget); in the work of Tong et al. [30], instead, a round denotes a time step in which either a node is selected, or the already active nodes influence some of their neighbors, and the goal is to select at most $k$ seeds that maximize the influence spread within $T$ rounds.
Moreover, running a campaign over multiple rounds can be very useful (maybe necessary) to the designer of the campaign in order to gain the necessary knowledge about the environment to address the desired goal. This approach has been widely studied under the lens of online learning [21]. Specifically, online learning algorithms for some submodular maximization processes (including the simplest form of our multi-round influence maximization problem) have been proposed [31, 12, 32]. In particular, these works introduced an approach, named Combinatorial Upper Confidence Bound (CUCB), that extended the well-known UCB bandit algorithm [4] with the aim of designing online learning algorithms with good guarantees for a wide class of optimization problems. However, such approach has been defined only for non-adaptive optimization frameworks, e.g., standard influence maximization problems in which the diffusion is not observed after each seed selection, but only at the end of the selection process. Gabillon et al. [20] presented CUCB algorithms for a class of adaptive optimization problems simpler than the one considered in this work, since, e.g., the seeding is subject to a limit on per-round budget and the underlying combinatorial structure is not sufficiently rich to represent influence maximization problems. Budget constraints without any limit on per-round quotas have been considered by Das et al. [16], that still achieves good guarantees, but considers an even simpler setting than the one by Gabillon et al [20], that in fact does not apply to influence maximization.
The analysis of optimization advantages of adaptive w.r.t. non-adaptive algorithms has been initiated by Dean et al [18, 17] on classical packing problems. Only recently, this kind of analysis has been applied to stochastic maximization problems, including single-round influence maximization and stochastic probing. In particular, Golovin and Krause [22] showed that whenever the objective function satisfies a property named adaptive submodularity (see Definition 2 below), a greedy algorithm achieves a $(1-\frac{1}{e})$-approximation of the optimal greedy algorithm. Chen and Peng [28, 11] also proved constant bound on the adaptivity gap for some influence maximization problems, i.e., by using a non-adaptive optimal algorithm we lose only a constant approximation factor w.r.t. the optimal adaptive algorithm. Some bounds on the approximation ratio of adaptive algorithms and on the adaptivity gap have been recently improved by D’Angelo et al [14, 15]. While our model and results are clearly inspired by these works, machineries defined therein cannot be directly applied to our setting since we are also interested in allocating the budget among multiple rounds.
Our Contribution.
In order to provide results that hold for both influence maximization and stochastic probing (and for variants or generalizations of them), we will first introduce the framework of Stochastic Budgeted Multi-round Submodular Maximization (SBMSm) problems. These problems are defined by a number $T$ of rounds; a ground set of $n$ items, whose state is drawn at each time step from an item-specific set according to some time-dependent distributions; a t; and a budget $B$ on the possible observations done within the time-horizon $T$. The problem is to choose which observations to make in each of the $T$ rounds, subject to the budget limit, in order to maximize the expected total value of the state function over the $T$ rounds.
This problem extends and generalizes both the influence maximization and stochastic probing problems (more details will be provided below): in the first case, the state of an item $v$ defines which nodes would be influenced if $v$ is selected as a seed, with the distribution of the state of $v$ depending on edges’ probabilities; an observation consists on choosing a node as a seed, running the influence spread, and evaluating the realized state by, e.g., counting the number or the total weight of influenced nodes. In the second case, the state of an item can be active or inactive, and we can observe the state of selected items only; the value of the submodular set function depends on the set of selected items which are active. Note that SBMSm not only generalizes influence maximization and stochastic probing problems to multiple round settings, but it also allows to model more realistic settings where states can change over time: e.g., in the influence maximization setting, we may allow edge probabilities or node weights to change in different rounds or be discounted over time.
We will start by characterizing when the problem can be optimally solved. In particular, we will show that whenever a certain single round optimization problem can be optimally solved in polynomial time, then there is a polynomial time dynamic programming algorithm that returns the optimal allocation of observations. Note that this optimal allocation is assumed to be very strong: at each time step, based on observations both in the past and in the current round, it decides if to spend a further unit of budget in the current round to make a new observation, or to stop the observations for the current round, and save budget for next rounds. Note that this will allow the optimum to react immediately to very negative events (e.g., the currently chosen seeds have influenced much less nodes than expected) or to very positive events (e.g., the currently chosen seeds have influenced much more nodes than expected). Hence, we find it is quite surprising that the optimality of the single round problem straightly extends to multiple rounds.
Unfortunately, the single-round problem turns out to usually be a hard problem to solve. Indeed, it can be easily reduced from the (deterministic) monotone submodular maximization problem, that is is known to be NP-hard [19], and no algorithm is known that approximate this problem in general instances within a factor better than $1-1/e$. Nevertheless, we show that there are non-trivial instances for which the problem can be effectively solved in polynomial time, essentially corresponding to the cases in which many observations will provide “sure” results, and essentially few of them are subject to randomness. A natural direction would be to extend our framework so that it would work also with similar but more general instances where, for example, one may compute an arbitrarily good approximation in polynomial time. However, we will show that within our dynamic programming framework this approximation ratio will explode with the number of rounds, and hence fails to provide good approximation guarantees.
For this reason, we need to resort to a different framework. Interestingly, we will show that a greedy polynomial time algorithm, that is arguably simpler than the previous one, provides a $(1/2-\epsilon)$-approximation of the optimal allocation for every instance. Hence, its approximation is very close to what happens in the much simpler single-round case [22].
Interestingly, this greedy approximation algorithm relies on a result that can be of independent interest. Indeed, we prove that in a SBMSs problem, a non-adaptive allocation of budget to rounds can be sufficient to guarantee a constant approximation of the adaptive optimal allocation of observations described above. Even more surprisingly, we can achieve such result by allocating the budget to rounds greedily, i.e., by assigning each unit of budget to the round in which it is expected to produce the largest expected value of the score function, given the amount of budget still available.
Our ($1/2-\epsilon$)-approximate algorithm then follows by composing this greedy allocation of budget, with the algorithm that chooses the items to observe up to the allocated budget in a greedy adaptive way, i.e., it chooses an item that is expected to achieve the maximum value of the state function, observes the realization of the state for this item, and then it chooses the next item among the ones that are expected to maximize the state function given the state realization for the first item, and so on. In other words, our result shows that the powerful optimal allocation of the observations described above, that is adaptive both in the allocation of budget to rounds and in the allocation of observations within rounds, can be well-approximated by an algorithm that is only partially adaptive, i.e., it is adaptive in the allocation of observations within rounds, but not in the allocation of budget to rounds. We point out that the effectiveness of a partially adaptive approach for multi-round adaptive optimization has been left as prominent research direction by [5], and to the best of our knowledge, our results are the first that consider such an approach and guarantee, at the same time, a good approximation to the optimal solution in polynomial time.
2 Model and Definitions
Before describing the framework of Stochastic Budgeted Multi-round Submodular Maximization, we introduce some preliminary notation, a general class of maximization problems over multiple rounds, and the main tools of adaptive optimization.
Preliminary Notation.
Given two integers $k_{1},k_{2}$, let $\llbracket k_{1},k_{2}\rrbracket:=\{k_{1},\ldots,k_{2}\}$ be the discrete interval from $k_{1}$ to $k_{2}$ (that is empty if $k_{1}>k_{2}$); furthermore, let $[k]$ denote the discrete interval $\llbracket 1,k\rrbracket$. Given an event $A$, let $\bm{1}_{A}$ denote the boolean variable that is equal to $1$ if $A$ is true/happens, and equal to 00 otherwise.
2.1 A Setting of Multi-round Maximization
We consider a setting $I:=(T,B,V,H,(\mathcal{P}_{t})_{t\in[T]},(f_{t})_{t\in[T]})$ whose parameters and connected properties are described as follows. $T\geq 2$ is a finite and integral horizon, where each $t\in[T]$ denotes the $t$-th round. $B\geq 0$ is a given budget. $V$ is a ground set of $n\geq 2$ items.
For any fixed $v\in V$, let $H(v)$ be a set of possible local states for item $v$, and let $H\subseteq\times_{v\in V}H(v)$ be the set of possible global states, that is, a global state $\eta\in H$ is a collection of local states $(\eta(v))_{v\in V}$ (with $\eta(v)\in H(v)$ for any $v\in V$). In the following, the global states may be simply called states.
For any round $t$, a state $\eta_{t}\in H$ is picked according to a probability distribution $\mathcal{P}_{t}$ over $H$, that is called (global) state distribution at round $t$; the state distributions $\mathcal{P}_{1},\ldots,\mathcal{P}_{T}$ are not generally independent. In the following, we will adopt bold letters to denote the random variables, and non-bold letters to denote their realization (e.g., $\bm{\eta}(v)$ is a random local-state, and $\eta(v)$ denotes a possible realization of $\bm{\eta}(v)$).
Finally, for any round $t\in[T]$, $f_{t}:2^{V}\times H\rightarrow{\mathbb{R}}_{\geq 0}$ is the state function at round $t$, that associates a non-negative real-value $f_{t}(S_{t},\eta_{t})$ with each subset of items $S_{t}\subseteq V$ and state $\eta_{t}$. Note that state functions are allowed to be different at each time step. Such a variability in both the state distribution and state function, immediately makes inefficient to allocate the same budget to each round: indeed, it is immediate to build a setting in which many rounds may lead to very poor outcomes, and only few of them may lead to a very large outcome, and hence all the budget allocated to the former rounds is wasted.
Let $\bar{f}:(2^{V}\times H)^{T}\rightarrow{\mathbb{R}}_{\geq 0}$ be the aggregated state function, that assigns the value $\bar{f}(\vec{S},\vec{\eta}):=\sum_{t=1}^{T}f_{t}(S_{t},\eta_{t})$ to each aggregated subset $\vec{S}:=(S_{1},\ldots,S_{T})$ and aggregated state $\vec{\eta}:=(\eta_{1},\ldots,\eta_{T})$, i.e., it returns the cumulative value of the state function $f$ over all rounds.
We consider then the problem of selecting an aggregated subset $\vec{S}=(S_{1},S_{2},\ldots,S_{T})$ such that $\bar{f}(\vec{S},\vec{\eta})$ is maximized subject to the budget constraint $\sum_{1}^{T}|S_{t}|\leq B$, for a fixed aggregated state $\vec{\eta}$ drawn from the aggregated state distributions $(\bm{\eta}_{1},\ldots,\bm{\eta}_{T})$. Unfortunately, we need to solve this problem without knowing the aggregated state, and we have access to the aggregated state distributions and functions only.
2.2 Maximization via Adaptive Policies
We here focus on algorithms for selecting the aggregated subset $\vec{S}$ achieving (an approximation of) the optimal value of $\bar{f}(\vec{S},\vec{\eta})$ that are adaptive, i.e., they select items over subsequent rounds, observe the partial states of the probed items and use these observations to guide the future choices in an adaptive fashion.
We next provide a formal definition for this adaptive behavior. Let us first introduce some further notation. For a given state $\eta_{t}$ at round $t$, we let $\langle S_{t},\eta_{t}\rangle=(\eta_{t}(v))_{v\in S_{t}}$ be the partial state observed when selecting the items in $S_{t}$ at round $t$. As the state $\eta_{t}$ is generally unknown, we will simply denote $\langle S_{t},\eta_{t}\rangle$ as $\langle S_{t}\rangle$. An aggregated partial state is a vector $\langle\vec{S}\rangle:=(\langle S_{1}\rangle,\ldots,\langle S_{T}\rangle)$ of partial states, one at each round, and collects the observations done at each round.
We look for a multi-round (adaptive) policy $\pi$, that on input the (unknown) aggregated state $\vec{\eta}$ returns an aggregated subset $\vec{S}:=\pi(\vec{\eta})$, according to the following procedure:
  1. 1.
    It initially starts from the first round and an empty aggregated set, and it initializes the adaptive strategy by setting $t=1$ (the actual round), $\vec{S}:=(S_{1},\ldots,S_{T}):=(\emptyset,\ldots,\emptyset)$ (the actual aggregated set), and $i=1$ (the actual step at round $t$) [^4] We highlight that we will use the term round to define the $T$ phases over which our optimization problem is defined. In each round, the adaptive policy can be in turn executed over multiple phases: these will be termed steps. Finally, we use the term time only with reference to computational time..
  2. 2.
    At each round $t$ and step $i$, either it (a) selects an item $v_{t,i}\in V$ not included in $S_{t}$ yet, or (b) it decides to stop and proceed to next round.
    • (a)
      If it selects an item $v_{t,i}$, it adds it to $S_{t}$ and sets $i\leftarrow i+1$, i.e., the policy is ready for the next item selection in the same round. Having selected the item $v_{t,i}$, it can observe the local state $\eta_{t}(v_{t,i})$. Hence, the policy knows the partial state $\langle S_{t},\eta_{t}\rangle$ and it can use such knowledge to select the (eventual) next item to be probed and included in $S_{t}$.
    • (b)
      If the policy decides to cease in selecting items in the current round, then it sets $t\leftarrow t+1$ and $i\leftarrow 1$.
  3. 3.
    Repeat step 2 until $t>T$ (i.e., the horizon has been exceeded) or $\sum_{t^{\prime}=1}^{t}|S_{t^{\prime}}|=B$ (i.e., all the budget has been used).
Among the possible multi-round policies, we are interested in the ones that return (an approximation of) the maximum $\sigma(\pi):=\mathbb{E}_{\vec{\bm{\eta}}\sim\mathcal{P}^{T}}\left[\bar{f}(\pi(\vec{\bm{\eta}}),\vec{\bm{\eta}})\right]$, (where $\pi(\vec{\bm{\eta}})$ is the aggregated set returned by $\pi$ and $\mathcal{P}^{T}$ denotes the product distribution ×$\times_{t\in[T]}\mathcal{P}_{t}$). Note that the optimal value achievable by a multi-round adaptive policy can be not so high as that obtained if we knew the input aggregated state, but it is the best we can hope by inferring the aggregated state via the adaptive probing of the items, based on the knowledge of the state distribution.
An interesting class of the above policies is that of semi-adaptive policies, in which the number of seeds that must be selected at each round (i.e., the budget per round) is determined non-adaptively in advance, and then the adaptive selection of items is performed within each round independently. Semi-adaptive policies are computationally simpler than general adaptive ones, but it might have worse performance guarantees. Actually, we will show that this worsening is limited for our problem.
Next, with a little abuse of notation, we will often use the notation $\pi(I,a_{1},\ldots)$ to denote a policy applied to $I$, where $a_{1},\ldots$ are further parameters defining the policy, and the (unknown) aggregated state $\vec{\eta}$ will be included as further input parameter.
2.3 Properties of the State Functions
In the following, we list some additional properties that the state functions must verify in our setting.
The first subset of properties will focus on the complexity of the state function. Specifically, we assume that each state function $f_{t}$ is representable and computable in polynomial time. Moreover, we also constraint the maximum value of a state function, as follows.
Definition 1.
A state function $f_{t}$ is polynomially bounded if there exists real values $\lambda,\Lambda>0$, where $\Lambda$ and $1/\lambda$ are polynomially bounded in the size of $I$, such that the following conditions hold: (ii) $f_{t}(\{v\}\cup S_{t},\eta_{t})-f_{t}(S_{t},\eta_{t})\leq\Lambda$ for any $v\in V$, $S_{t}\subseteq V$ and state $\eta_{t}$; (ii) there exist $t\in[T]$ and $v\in V$ such that $\lambda\leq\mathbb{E}_{\bm{\eta}_{t}\mathcal{P}_{t}}[f_{t}(\{v\},\bm{\eta}_{t})]$.
The second subset of properties instead extends the typical feature of monotonicity and submodularity to our multi-round adaptive setting. Specifically, we assume that, for any fixed state $\eta_{t}$, the function $f_{t}$ satisfies the following monotonicity property: $f_{t}(S_{t},\eta_{t})\leq f_{t}(S_{t}^{\prime},\eta_{t})$ if $S_{t}\subseteq S_{t}^{\prime}$.
We further assume that state functions satisfy a certain property connected with the notion of submodularity. Given two partial states $\langle S_{t}\rangle:=\langle S_{t},\eta_{t}\rangle$ and $\langle S^{\prime}_{t}\rangle:=\langle S^{\prime}_{t},\eta^{\prime}_{t}\rangle$, we say that $\langle S_{t}\rangle$ is a sub-state of $\langle S^{\prime}_{t}\rangle$ (equivalently, we say $\langle S_{t}\rangle\preceq\langle S^{\prime}_{t}\rangle$) if $S_{t}\subseteq S^{\prime}_{t}$, and $\eta_{t}(e)=\eta^{\prime}_{t}(e)$ for any $e\in S_{t}$, that is, the local states observed in $\langle S_{t}\rangle$ can be also observed in $\langle S^{\prime}_{t}\rangle$. Given a partial state $\langle S_{t}\rangle$ at round $t$ and $v\in V$, let $\Delta_{t}(v\mid\langle S_{t}\rangle):=\mathbb{E}_{\bm{\eta}_{t}\sim\mathcal{P}}\left[f_{t}(S_{t}\cup\{v\},\bm{\eta}_{t})-f(S_{t},\bm{\eta}_{t})|\langle S_{t}\rangle\preceq\bm{\eta}_{t}\right]$ i.e., the expected increment of $f_{t}$ obtained when adding item $v$ to $S_{t}$, conditioned by the observation of partial state $\langle S_{t}\rangle$.
We then have the following definition, that adapts the notion of adaptive submodularity introduced by [22] to our setting.
Definition 2 ([22]).
A state function $f_{t}:2^{V}\rightarrow{\mathbb{R}}_{\geq 0}$ is adaptive submodular if, for all $\langle S_{t}\rangle\preceq\langle S^{\prime}_{t}\rangle$ and $v\in V$ we get $\Delta_{t}(v\mid\langle S_{t}\rangle)\geq\Delta_{t}(v\mid\langle S^{\prime}_{t}\rangle)$.
Henceforth, we assume that state functions satisfy all the properties described above.
2.4 The Stochastic Budgeted Multi-round Submodular Maximization Problem
We denote the setting described above as Stochastic Budgeted Multi-round Submodular Maximization Problem (SBMSm), that looks for a multi-round policy $\pi$ maximizing $\sigma(\pi)=\mathbb{E}_{\vec{\bm{\eta}}\sim\mathcal{P}^{T}}\left[\bar{f}(\pi(\vec{\bm{\eta}}),\vec{\bm{\eta}})\right]$ on an input instance $I=(T,B,V,H,(\mathcal{P}_{t})_{t\in[T]},(f_{t})_{t\in[T]})$.
In particular, we will focus on policies with guaranteed approximation, defined as follows. Given an instance $I=(T,B,V,H,(\mathcal{P}_{t})_{t\in[T]},(f_{t})_{t\in[T]})$ of SBMSm, $\alpha\in[0,1]$, and a multi-round policy $\pi$ for $I$, we say that $\pi$ is $\alpha$-approximate, if $\sigma(\pi)\geq\alpha\cdot\sigma(\pi^{*})-\beta$, where $\pi^{*}$ denotes an optimal multi-round policy for $I$.
In the remainder of the paper, we will implicitly assume that $B<nT$, where $B$ is the budget of the input instance. Indeed, if $B\geq nT$, we could select a set $V$ at each round so that the state function is trivially maximized at each round. Furthermore, the assumption $B<nT$ implies that the budget $B$ is polynomial in the size of the input instance.
Before providing our results, we will provide a description on how to apply our framework to multi-round and adaptive variants of influence maximization and stochastic probing.
2.5 Applications
2.5.1 Influence Maximization
The Influence Maximization problem has been extensively studied in the last years under different models of information diffusion. Here we focus on the Independent Cascade diffusion model but a similar approach can be used to model the Influence Maximization under different diffusion models. Before to show how this problem can be fitted in the class of SBMSm, let us first define the problem of Influence Maximization under the Information Cascade model and introduce its adaptive and multi-round variants.
Definition.
The Independent Cascade diffusion model (IC) [26], given a graph $G=(V,E)$, a probability $p\colon E\rightarrow[0,1]$ defined on edges, an obstruction $F\subseteq V$, and a set of seeds $S_{0}\subseteq V\setminus F$, returns a set $A$ of active nodes computed as follows: for each round $t\geq 0$ such that $S_{t}\neq\emptyset$, $S_{t}$ is added to $A$, and then for each edge $(u,v)\in E$ such that $u\in S_{t}$ and $v\notin A\cup F$, add $v$ to $S_{t+1}$ with probability $p(u,v)$.
In the problem of Influence Maximization under Independent Cascade diffusion model (IM-IC) [26] we are given a graph $G=(V,E)$, a weight function $w\colon V\rightarrow\mathbb{R}_{+}$ defined on vertices, a probability $p\colon E\rightarrow[0,1]$ defined on edges, and a budget $B$, and we are asked to select the set $S^{*}$ of at most $B$ vertices that maximizes the expected sum of weights of active nodes returned by the Independent Cascade diffusion model on input $G$, $p$, an empty obstruction and $S$, i.e., $S^{*}=\arg\max_{S\subseteq V\colon|S|\leq B}\mathbb{E}_{p}\left[\sum_{v\in IC(G,p,\emptyset,S)}w(v)\right]$.
Recently, an adaptive variant of the problem has been proposed where seeds are introduced one at a time, using an adaptive policy. To define the adaptive variant of the problem we need first to define the concept of adaptive policy. Given a graph $G=(V,E)$, a probability $p\colon E\rightarrow[0,1]$ defined on edges, and a budget $B$, an adaptive policy $\pi$ for IM-IC works as follows: it selects a first vertex $v_{1}$, observes the set $A_{1}$ of vertices activated by seed $v$, i.e., $A_{1}=IC(G,p,\emptyset,\{v_{1}\})$, then selects a new vertex $v_{2}\in V\setminus A_{1}$, observes the set $A_{2}$ of newly activated vertices, i.e., $A_{2}=IC(G,p,A_{1},\{v_{2}\})$, then selects a new vertex $v_{3}\in V\setminus(A_{1}\cup A_{2})$, observes the set $A_{3}$ of newly activated vertices, i.e., $A_{3}=IC(G,p,A_{1}\cup A_{2},\{v_{3}\})$, and so on until either there is no vertex that can be selected or $B$ vertices have been already selected. The set of nodes activated through an adaptive policy $\pi$ is $A(\pi)=\bigcup_{i\in[B]}A_{i}$.
In the problem of Adaptive Influence Maximization under Independent Cascade diffusion model [22] we are given a graph $G=(V,E)$, a weight function $w\colon V\rightarrow\mathbb{R}_{+}$ defined on vertices, a probability $p\colon E\rightarrow[0,1]$ defined on edges, and a budget $B$, and we are asked to select the adaptive policy $\pi^{*}$ that maximizes the expected sum of weights of activated nodes, i.e., $\pi^{*}=\arg\max_{\pi}\mathbb{E}_{p}\left[\sum_{v\in A(\pi)}w(v)\right]$.
In this paper we are interested to multi-round settings. Hence, we need to extend previous concepts to the multi-round setting. Given a graph $G=(V,E)$, a probability $p_{t}\colon E\rightarrow[0,1]$ defined on edges (for each round $t$), a budget $B$, and a time horizon $T$, a multi-round adaptive policy $\pi$ for IM-IC works as follows: at each round $t\in[T]$, let $S^{t}(\pi)$ be the set of seeds selected by $\pi$ in round $t$ and $A^{t}_{v}(\pi)$ be the set of nodes activated by $v\in S^{t}(\pi)$ (we will next omit the reference to $\pi$ when clear from the context). Note that both $S^{t}$ and $A^{t}_{v}$ are initially empty. Then, the policy at each round $t\in[T]$, if the total number of selected vertices is lower than $B$,
  1. 1.
    decides whether to choose another seed in the current round or to move to the next round (note that if all vertices have been activated in the current round, then the policy must necessarily choose to move to the next round);
  2. 2.
    in the first case, it selects a vertex $v$ that has not been yet activated in the current round, i.e., $v\notin A^{t}_{S^{t}}=\bigcup_{u\in S^{t}}A^{t}_{u}$, observes the set $A^{t}_{v}$ of newly activated vertices, i.e., $A^{t}_{v}=IC(G,p,A^{t}_{S^{t}},\{v\})$, adds $v$ to $S^{t}$, and repeat from step 1.
Finally, in the problem of Multi-Round Adaptive Influence Maximization under Independent Cascade diffusion model (MR-IM-IC) we are given a graph $G=(V,E)$, a time horizon $T$, a weight function $w_{t}\colon V\rightarrow\mathbb{R}_{+}$ defined on vertices for every $t\in[T]$, a probability $p_{t}\colon E\rightarrow[0,1]$ defined on edges for each round $t\in[T]$, and a budget $B$, and we are asked to select the multi-round adaptive policy $\pi^{*}$ that maximizes the expected sum of weights of activated nodes over $t$ rounds, i.e., $\pi^{*}=\arg\max_{\pi}\varsigma(\pi)$ where $\varsigma(\pi)=\mathbb{E}_{(p_{t})_{t\in[T]}}\left[\sum_{t\in[T]}\sum_{v\in A^{t}_{S^{t}}(\pi)}w_{t}(v)\right]$.
We stress that our definition allows that both vertices’ weights and edge probabilities may change from one round to the next: this can be useful to implement a variable discount factor (e.g., we care more about the number of people influenced in the last phases of the campaign) or to implement a dynamic environment in which the members of a population are known to appear and disappear over time, or the relevance of population members changes over time (e.g., in liquid democracy settings [6] at the beginning we are more interested in influencing a large portion of the population for manipulating the choice of delegates, while later in the campaign we are more concerned in influencing those members that have more chances to be chosen as delegates, in order to manipulate the outcome of their voting process).
Mapping between IM and SBMSm.
We start by showing how an instance of the MR-IM-IC problem can be mapped to an instance of SBSMm. Let $M=(T,B,V,E,(p_{t})_{t\in[T]},(w_{t})_{t\in[T]})$ be an instance of MR-IM-IC and let $I=(T,B,V,H,(\mathcal{P}_{t})_{t\in[T]},(f_{t})_{t\in[T]})$ be an instance of SBSMm. The mapping works as follows:
  • $T$ and $B$ are exactly the same as in $M$ (hence, we use exactly the same name);
  • the set $V$ of items is the set of nodes in the social network;
  • the set $H(v)$ of possible local states of $v$ corresponds to the possible sets of nodes influenced by an independent cascade starting from node $v$, i.e., $H(v)=\{X\subseteq V\mid v\in X\}$;
  • the probability $\mathcal{P}_{t}(\eta_{t})$ is the probability that an independent cascade influences exactly the $\eta_{t}(v)\in H(v)$ nodes when started from $v$ for every $v\in V$. Specifically, we can define $\mathcal{P}_{t}$ as follows: for each spanning subgraph $G_{E^{\prime}}$ of $G$, i.e., $G_{L}=(V,E^{\prime})$ with $E^{\prime}\subseteq E$, we set $P_{t}(G_{E^{\prime}})=\prod_{e\in E\cap E^{\prime}}p_{t}(e)\prod_{e\in E\setminus E^{\prime}}(1-p_{t}(e))$, i.e. $G_{E^{\prime}}$ denotes one of the possible live graphs and $P(G_{E^{\prime}})$ its probability; for each $v$ and each $E^{\prime}\subseteq E$, we denote as $R_{E^{\prime}}(v)$ the set of nodes reachable from $v$ in $G_{E^{\prime}}$; for a state $\eta_{t}$, we set $\Xi(\eta_{t})=\{E^{\prime}\subseteq E\mid\eta_{t}(v)=R_{E}^{\prime}(v)\forall v\in V\}$, i.e., $\Xi(\eta_{t})$ is the set of live graph that are compatible with the state of all vertices described in $\eta_{t}$; we say that $\eta_{t}$ is feasible if $|\Xi(\eta_{t})|\geq 1$; then, $\mathcal{P}_{t}(\eta_{t})=0$ if $\eta_{t}$ is infeasible, and $\mathcal{P}_{t}(\eta_{t})=\sum_{E^{\prime}\in\Xi(\eta_{t})}P_{t}(G_{E^{\prime}})$;
  • the state function $f_{t}$ is defined as $f_{t}(S,\eta_{t})=\sum_{v\in\bigcup_{s\in S}\eta_{t}(s)}w_{t}(v)$.
According to this mapping between $M$ and $I$, and the obvious mapping between a multi-round policy $\pi$ for IM-IC and a multi-round policy $\pi^{\prime}$ for SBMSm, it is immediate to check that $\varsigma(\pi)=\sigma(\pi^{\prime})$, and thus finding the policy maximizing the latter is equivalent to find a policy maximizing the former.
Properties of $f_{t}$.
In order to establish that $I$ is a valid instance for SBMSm we need also to prove that $f_{t}$ is polynomially computable and representable, it is polynomially bounded, and it satisfies monotonicity and adaptive submodularity.
Polynomial representation and computation immediately follows from the definition above. As for polynomial boundedness, this property is immediately satisfied if there exist constants $c,C>0$ such that $w_{t}(v)\leq C$ for every $v$ and every $t$, and there exist a round $t$ and a node $v$ such that $c\leq w_{t}(v)$. In this case we have $\Lambda\leq Cn$ and $\lambda=c$.
Monotonicity is immediately satisfied since the information diffusion cannot decrease if the set of observed edges expands. As for the adaptive submodularity, it is not hard to check that this condition is satisfied: intuitively, for each realization of the live graph, as described above, the increment of $f_{t}$ due to $v$ is given by nodes reachable from $v$ in this live graph, but not reachable by previously selected seeds; and it is clearly impossible to expand this subset with more previously selected seeds.
2.5.2 Stochastic Probing
Definition.
In general stochastic probing problems, we have $n$ items, each one associated with a random state independent from the others, and the goal is to choose a feasible subset of items maximizing the value of a certain function depending on the states of the selected items. Several classes of stochastic probing problems arise from particular choices of the states and the function to be maximized.
The stochastic probing problem we consider here, denoted as binary submodular stochastic probing (BSS), is defined as follows: We have a set $V$ of $n$ items, and a non-negative, monotone and submodular set function $g:2^{V}\rightarrow{\mathbb{R}}_{\geq 0}$ that we aim at maximizing. Each item $v$ has a binary state $\phi_{v}$ in the set $\{active,inactive\}$, the probability that $v$ is a active is $p(v)$, and the distributions associated to all items are independent. Given a subset $S\subseteq V$ and a realization $\phi=(\phi(v))_{v\in V}$ of the states, let $act(S,\phi)\subseteq S$ denote the subset of active items in $S$. For a given budget $B\geq 0$, the goal is to find a subset $S^{*}\subseteq V$ of at most $B$ items that maximizes $\mathbb{E}_{\phi}[g(act(S^{*},\phi))]$ (i.e., the expected value of $g$ applied to the selected items which are active).
The multi-round adaptive variant of this problem, that we call multi-round binary submodular stochastic probing (MR-BSS) assumes that the probabilities and the submodular set functions are time specific (i.e., we have a probability distribution $p_{t}$ and a function $g_{t}$ for any round $t$), the budget (i.e., the number of items we can select) can be adaptively allocated over all possible rounds, and the goal is to maximize the sum of the values achieved by all functions $g_{t}$. In the adaptive variant of the problem, one considers adaptive multi-round policies that adaptively select the items at each round. Furthermore, we assume that there are two quantities $\lambda,\Lambda>0$ that are polynomial in the input instance such that: (i) the increment caused to $g_{t}$, when probing a new item $v$, is upper bounded by $\Lambda$; (ii) there exists one round $t$ and an item $v$ such that the expected value of $g_{t}$ when probing item $v$ at round $t$ is at least $\lambda$.
Mapping between MR-BSS and SBMSm.
Let $M=(T,B,V,(p_{t})_{t\in[T]},(g_{t})_{t\in[T]})$ be an instance of the MR-BSS problem, and let us now show how this instance can be modelled as an SBMSm instance $I=(T,B,V,H,(\mathcal{P}_{t})_{t\in[T]},(f_{t})_{t\in T})$: $T$, $B$, and $V$ are exactly the same as in $M$; $H(v)=\{0,1\}$, with 00 meaning that item $v$ is inactive, and $1$ meaning that it is active, and thus $H=\{0,1\}^{n}$; $\mathcal{P}_{t}(v)=p_{t}(v)$; and $f_{t}(S,\eta_{t})=g_{t}(\{v\in S\colon\eta_{t}(v)=1\})$, i.e., we evaluates among the set $S$ of probed items, only the ones that happens of being active.
According to this mapping between $M$ and $I$, we have an obvious mapping between a multi-round policy $\pi$ for MR-BSS and a multi-round policy $\pi^{\prime}$ for SBMSm, thus finding the policy maximizing the latter is equivalent to find a policy maximizing the former.
Properties of $f_{t}$.
All desired properties of $f_{t}$ follows if $g_{t}$ also enjoys the same properties by definition.
3 Optimal Algorithm
In this section we first describe a single-round optimization problem of interest, and how its solution is used to provide a dynamic programming algorithm that will return the optimal policy in polynomial time as long as the single-round optimization problem can be efficiently addressed. Then, we will discuss a few of cases in which the single-round optimization problem can be efficiently addressed. Finally, we will highlight the limits of this dynamic programming framework, by discussing when it can be extended to provide a good approximation of the optimal policy, even if the single-round optimization problem can only be approximated.
3.1 A Dynamic Programming Algorithm based on Single-round Optimization
Let us fix a round $t$ and a budget $b$ available. Given $j\in\llbracket 0,b\rrbracket$, let $R(t+1,j)$ denote the optimal expected value of $\sum_{t^{\prime}=t+1}^{T}f_{t^{\prime}}$ achievable in rounds $t+1,\ldots,T$ by a multi-round policy $\pi_{\geq t+1}$ restricted to such rounds only; we observe that $R(t+1,j)=0$ for $t=T$. We assume that $R(t+1,k)$ has been already determined for any $k\in\llbracket 0,b\rrbracket$. The single-round optimization problem at round $t$ asks to find a policy $\pi_{t}$ at round $t$ that adaptively computes the budget $\bm{i}\in\llbracket 0,b\rrbracket$ to be allocated at round $t$ in such a way that the expected value
$\mathbb{E}_{\bm{\eta}_{t}\sim\mathcal{P}_{t}}[f_{t}(\pi_{t}(\bm{\eta}_{t}),\bm{\eta}_{t})+R(t+1,b-\bm{i})]$ (1)
of state function $f_{t}$ plus $R(t+1,b-\bm{i})$ is maximized, where the random budget $\bm{i}$ used at round $t$ depends on the chosen policy $\pi_{t}$ and the state $\bm{\eta}_{t}$.
We observe that the optimal value of (1) coincides with $R(t,b)$, that is, the optimal expected value of $\sum_{t^{\prime}=t}^{T}f_{t^{\prime}}$ achievable in rounds $t,\ldots,T$. Thus, the following recursive relation holds:
$R(t,b)=\begin{cases}0,&\text{ if }t=T+1\\ \max_{\pi_{t}}\mathbb{E}_{\bm{\eta}_{t}\sim\mathcal{P}_{t}}[f_{t}(\pi_{t}(\bm{\eta}_{t}),\bm{\eta}_{t})+R(t+1,b-\bm{i})],&\text{ if }t\in\llbracket 0,T\rrbracket.\end{cases}$ (2)
Suppose that there is an algorithm for the single-round optimization problem described above, that is, by resorting to the recursive relation (2), we can efficiently compute $R(t,b)$ as a function of $R(t+1,j)$ for any $j\in\llbracket 0,b\rrbracket$, together with the optimal single-round policy. Such algorithm can be used as a sub-routine for a more general dynamic programming algorithm that computes an optimal policy for the general SBMSm problem.
Specifically, the dynamic programming algorithm first sets $R(T+1,b)=0$ for every $b\in\llbracket 0,B\rrbracket$. By resorting to the recursive relation (2), we can compute $R(T,b)$ for any $b\in\llbracket 0,B\rrbracket$, as a function of the values $R(T+1,j)$ with $j\in\llbracket 0,b\rrbracket$. Iteratively, going from $t=T-1$ to $t=1$, we can compute $R(t,b)$ for any $b\in\llbracket 0,B\rrbracket$, as a function of the values $R(t+1,j)$ with $j\in\llbracket 0,b\rrbracket$. Finally, we observe that $R(1,B)$ is the optimal value for the SBMSm problem. Thus, starting from round $t=1$ and budget $B$, we can reconstruct the optimal multi-round policy achieving the optimal value (this can be done by resorting to standard dynamic programming techniques).
3.2 How to Solve Single-round Optimization Efficiently?
In the above dynamic-programming algorithm, we assumed the existence of a sub-routine that solves the single-round optimization problem, i.e., that allows to compute recurrence (2). Unfortunately, the considered single-round problem can be easily reduced from the submodular monotone optimization problem, that is NP-hard [19]. To deal with the intractability of the single-round optimization problem, we will describe an algorithm that is in general inefficient, but that can be run in polynomial time in a few cases.
The Single-round Optimization Problem as a Game.
To compute a single-round policy $\pi_{t}$ that computes $R(t,b)$ for any round $t$ and budget $b$, it is useful to describe the single-round optimization problem at round $t$ in terms of a suitable two-player extensive form game between two players [27]: (i) the policy $\pi_{t}$, that decides, at each step, either to interrupt the game and collecting a reward or to continue by selecting another item; (ii) the nature $\mathcal{N}$, that reveals a partial state after each item selection.
Specifically, the game is modeled as a directed rooted tree $\mathcal{T}_{t}$ of depth $12n+$ 1. The policy $\pi_{t}$ will play at all nodes at even depth, whereas the nature $\mathcal{N}$ plays at all nodes at odd depth. We will denote with $\mathcal{T}_{t}^{o}$ the set of nodes in which $\pi_{t}$ plays, and with $\mathcal{T}_{t}^{e}$ the set of nodes in which ${\cal N}$ plays. Each leaf of the tree corresponds to an outcome of the game, in which the policy $\pi_{t}$ collects a reward depending on the reached leaf.
In the following, we define the tree $\mathcal{T}_{t}$, so as the game between $\pi_{t}$ and ${\cal N}$, starting from the nodes at depth $d\in\{0,1\}$, then at depth $d\in\{2,3\}$, and iteratively at depth $d\in\{2(i-1),2i-1\}$, for any $i\in[n+1]$.
Case $d\in\{0,1\}$: The root $u^{*}=u_{0}$ of $\mathcal{T}_{t}$, that belongs to $\mathcal{T}_{t}^{o}$ (as its depth is 00), is the initial configuration of the game. The root has $n+1$ outgoing edges, where each edge $(u_{0},u_{1})$ has a label $a=L(u_{0},u_{1})$ corresponding to the action $a$ that $\pi_{t}$ can perform at step $1$. In particular, either the action/label $a$ is a string STOP that will denote the termination of the single-round policy, or $a$ is an item $v_{1}\in V$ that $\pi_{t}$ can select at step $1$. Given the action $a$ chosen by $\pi_{t}$ at step $1$, the new configuration of the game is obtained by following the corresponding labeled edge (i.e., the edge $(u_{0},u_{1})$ such that $a=L(u_{0},u_{1})$), and it corresponds to a node $u_{1}$ at depth $1$. If $a=$STOP then the reached node $u_{1}$ is a leaf and the game ends. In such a case, the reward of $\pi_{t}$ is $R(t+1,b)$, as $\pi_{t}$ did not select any node and the whole budget $b$ can be used in rounds $t+1,\ldots,T$. Otherwise, the reached node $u_{1}$ has as many outgoing edges as partial states $\langle\{v_{1}\}\rangle$ that can be observed after selecting the item $v_{1}$, and each outgoing edge $(u_{1},u_{2})$ is labeled with a distinct partial state. At this point the nature $\mathcal{N}$ picks a partial state $\langle\{v_{1}\}\rangle$, drawn according to the probability distribution induced by $\mathcal{P}_{t}$ on partial states of type $\langle\{v_{1}\}\rangle$ (i.e., observable after selecting $v_{1}$). Then, the new configuration of the game is obtained by following the outgoing edge labeled with the picked partial state (i.e., the edge $(u_{1},u_{2})$ such that $\langle\{v_{1}\}\rangle=L(u_{1},u_{2})$), and it corresponds to a node $u_{2}$ at depth $2$, that belongs to $\mathcal{T}_{t}^{o}$.
Case $d\in\{2,3\}$: Now, $u_{2}$ has $n$ outgoing edges, again labeled with the actions $a$ that $\pi_{t}$ can perform at step $2$. In particular, either $a$=STOP, or $a$ is equal to an item $v_{2}\in V\setminus\{v_{1}\}$ that $\pi_{t}$ can select at step $2$. Given the action $a$ chosen by $\pi_{t}$ at step $2$, the new configuration of the game is obtained by following the corresponding labeled edge (i.e., the edge $(u_{2},u_{3})$ such that $a=L(u_{2},u_{3})$), and it corresponds to a node $u_{3}$ at depth $3$. If $a=$STOP then $u_{3}$ is a leaf, the game ends and the reward for $\pi_{t}$ is the expected value of $f_{t}$ constrained by the observation of partial state $\langle\{v_{1}\}\rangle$, plus the optimal expected value $R(t+1,b-1)$ achievable in the subsequent rounds $t+1,\ldots,T$ by using the residual budget (i.e., $b-1$ units). Otherwise, the reached node $u_{3}$ has as many outgoing edges $(u_{3},u_{4})$ as partial states $\langle\{v_{1},v_{2}\}\rangle$ that can be observed after selecting $v_{2}$ (constrained by the previous observation of $\langle\{v_{1}\}\rangle$), and each outgoing edge is labeled with a distinct partial state. At this point the nature $\mathcal{N}$ picks a partial state $\langle\{v_{1},v_{2}\}\rangle$, drawn according to the probability distribution induced by $\mathcal{P}_{t}$ on partial states of type $\langle\{v_{1},v_{2}\}\rangle$, conditioned by the observation of $\langle v_{1}\rangle$. The new configuration of the game is obtained by following the outgoing edge labeled with the picked partial state, thus reaching a node $u_{4}$ at depth $4$, that belongs to $\mathcal{T}_{t}^{o}$.
Case $d\in\{2(i-1),2i-1\}$: Iteratively, for any $i\in[n+1]$, each node $u_{d}\in\mathcal{T}_{t}$ at depth $d=2(i-1)$ corresponds to a possible outcome of policy $\pi_{t}$ obtained at the end of step $i-1$, where $i-1$ items $v_{1},\ldots,v_{i-1}$ have been already selected and a partial state $\langle\{v_{1},\ldots,v_{i-1}\}\rangle$ has been observed. $u_{d}$ has $n-i+2$ outgoing edges $(u_{d},u_{d+1})$, each one labeled with the action $a=L(u_{d},u_{d+1})$ that $\pi_{t}$ can performs at step $i$: either $a$=STOP, or $a$ is equal to an item $v_{i}\in V\setminus\{v_{1},\ldots,v_{i-1}\}$ that $\pi_{t}$ can select at step $i$; we observe that, for $i=n+1$, there is a unique outgoing edge labeled with action $a=$STOP. Again, given the action $a$ chosen by $\pi_{t}$ at step $i$, the new configuration of the game is obtained by following the corresponding labeled edge and reaching a node $u_{d+1}$ at depth $d+1$. If $a=$STOP, then $u_{d+1}$ is a leaf, the game ends and the profit for $\pi_{t}$ is the expected value of $f_{t}$ constrained by the observation of partial state $\langle\{v_{1},\ldots,v_{i-1}\}\rangle$, plus the optimal expected value $R(t+1,b-i+1)$ achievable in the subsequent rounds by using the residual budget (i.e., $b-i+1$ units). Otherwise, $u_{d+1}$ has as many outgoing edges as partial states $\langle\{v_{1},\ldots,v_{i}\}\rangle$ that can be observed after selecting $v_{i}$ (constrained by the previous observation of $\langle\{v_{1},\ldots,v_{i-1}\}\rangle$), and each outgoing edge is labeled with a distinct partial state. At this point the nature $\mathcal{N}$ picks a partial state $\langle\{v_{1},\ldots,v_{i}\}\rangle$ drawn according to the probability distribution induced by $\mathcal{P}_{t}$ on partial states of type $\langle\{v_{1},\ldots,v_{i}\}\rangle$, conditioned by the observation of $\langle\{v_{1},\ldots,v_{i-1}\}\rangle$. The new configuration of the game is obtained by following the outgoing edge labeled with the picked partial state, and it corresponds to a node $u_{d+1}$ at depth $d+1$.
From the Game to the Optimal Solution.
Given a node $u\in\mathcal{T}_{t}^{o}$, let $\langle S(u)\rangle$ denote the partial state observed when the configuration of the game is determined by $u$. We have that the above surjective map between nodes of $\mathcal{T}_{t}^{o}$ and partial states defines an equivalence between the tree-based game and the problem of finding a policy $\pi_{t}$ that maximizes (1). Indeed, let $a^{*}(u)$ denote the action that maximizes the expected value of the reward collected by $\pi_{t}$, assuming that the actual configuration of the game is $u$, and let $R(u)$ denote the resulting optimal expected value. We have that $a^{*}(u)$ is also the action that $\pi_{t}$ must follow for any partial state $\langle S(u)\rangle$ in order to maximize (1). Then, denoting the root of $\mathcal{T}_{t}$ by $u^{*}$, we have that $R(u^{*})$ coincides with $R(t,b)$, that is the optimal value of the single-round optimization problem at round $t$ with budget $b$.
We observe that, by following standard approaches to find the optimal strategy of a player in extensive-form games [27], $a^{*}(u)$ and $R(u)$ can be computed for any $u\in\mathcal{T}_{t}^{o}$ by recursively exploring the tree $\mathcal{T}$. By the equivalence bewteen the tree-based game and the single-round optimization problem, the optimal strategy for the former can be easily turned into an optimal solution for the latter.
When the Single-round Problem can be Efficiently Solved?
By using the above tree-based algorithmic procedure, the single-round optimization problem can be solved efficiently whenever the game tree ${\cal T}_{t}$ contains a number of nodes that is polynomial in the size of the input instance of the general SBMSm problem. To this aim, in the proposition below, we relate the size of ${\cal T}_{t}$ to the number of items and states, and we next show that there may be practical settings in which this is not rare.
Proposition 1.
${\cal T}_{t}$ has $O(2^{n}n!|H|)$ nodes.
Proof.
Let $z$ denote the number of leaves of ${\cal T}_{t}$, and let $|{\cal T}_{t}|$ denote the number of nodes of ${\cal T}_{t}$. We first show that
$|{\cal T}_{t}|\in O(z).$ (3)
By exploiting the structure of $\mathcal{T}_{t}$, we have that each non-leaf node $u\in\mathcal{T}_{t}^{o}$ has exactly one child node that is a leaf. Thus, the leaves of $\mathcal{T}$ are at least as numerous as the nodes of $\mathcal{T}_{t}^{o}$, that is, $|\mathcal{T}_{t}^{o}|\leq z$. As the father of each node of ${\cal T}^{e}_{t}$ belongs to ${\cal T}^{o}_{t}$, we necessarily have $|{\cal T}^{e}_{t}|\geq|{\cal T}^{o}_{t}|$. By putting the above inequalities together, we get $|{\cal T}_{t}|=|{\cal T}^{o}_{t}|+|{\cal T}^{e}_{t}|\leq 2|{\cal T}^{e}_{t}|\leq 2z$, and this shows (3).
Now, we show that
$z\in O(2^{n}n!|H|).$ (4)
Let $G$ be the function that maps each sequence $s$ of distinct items in $V$ and partial state state $\eta_{t}\in H$, to the leaf $u=G(s,\eta_{t})$ of $\mathcal{T}_{t}$ associated with the outcome of $\pi_{t}$ obtained when the realized state is $\eta_{t}$ and $s$ is the sequence of items selected by $\pi_{t}$ at round $t$. We observe that, by the structure of $\mathcal{T}_{t}$, function $G$ is surjective. Thus, as the number of sequences of distinct items in $V$ is at most $2^{n}n!$ (at most $n!$ orderings for each of the $2^{n}$ subsets of $V$), the number of leaves is at most $2^{n}n!|H|$, and this shows (4). By putting (3) and (4) together, the claim of the proposition follows. ∎
If $n!$ and $|H|$ are polynomial in the input instance of SBMSm, then $2^{n}n!|H|$ is polynomial, too. In such a case, by Proposition 1, the number of nodes of ${\cal T}_{t}$ is polynomial. Thus, the tree-based algorithm for the single-round optimization problem can be run in polynomial time, and consequently so can then the dynamic programming algorithm for SBMSm.
Note that $n!$ is polynomial if, for instance, the number $n$ of items is very small, e.g., constant. Such a limitation is not so rare in several scenarios of interest, as it often occurs that in large populations only few items are really relevant (e.g., among the many friends on social networks only few of them are able to affect our opinions, or among the many potential job candidates only few effectively satisfy all the requirements at which one may be interested in).
The limitation on the number of states $|H|$ occurs in many settings where the number of observable events is not so high. For instance, it could be the case of certain influence maximization problems where edge probabilities are correlated, and depends on a few possible stochastic events. Furthermore, it often occurs that, if the number of items is constant, the number of states $|H|$ is constant, too. For instance, in both influence maximization and stochastic probing, we have that the number of states, even if it grows exponentially in the number of items, it depends on such quantity only. Therefore, if the number of items is constant, the number of states will be constant, too.
What Happens if the Single-round Problem cannot be Efficiently Solved?
While the single-round optimization problem is known to be NP-hard, it would be still possible to efficiently find arbitrarily good approximation of their optimal solution in many settings. Hence, it would be natural to replace in our dynamic programming framework the optimal solution to the single-round optimization problem with its approximation. However, it is not hard to observe that this causes an explosion on the approximation guarantee: indeed, an $\alpha$- approximation of the optimal policy at round $t$ evaluated on a game tree in which the values of $R(t+1,b)$ for the different choices of $b$ are only $\beta$-approximations of the maximum expected revenue achievable in round $t+1,\ldots,T$ with budget $b$, may be not able to return a policy guaranteeing an approximation better than $\alpha\cdot\beta$ of the the maximum expected revenue achievable in round $t+1,\ldots,T$ with the given budget. Hence, by running a $\rho$-approximation algorithm for all $T$ steps, may only allows to compute a $\rho^{T}$-approximation of the optimal multi-round adaptive policy, that is clearly an unsatisfying approximation guarantee for every constant $\rho$.
While it would be possible to fix our dynamic programming framework in order to do not cause such an explosion in the approximation guarantee, and to provide an $\left(\frac{1}{2}-\varepsilon\right)$-approximation to the optimal multi-round policy, we next prove that the same bound can be actually achieved by a simpler and less computationally expensive greedy algorithm.
4 Semi-Adaptive Approximation Algorithms
We now consider a semi-adaptive policy based on the greedy approach, that works as follows: it first determines how to distribute the budget $B$ over all rounds by computing a budget vector $\vec{b}=(b_{1},\ldots,b_{T})$, so that $b_{t}$ is the budget assigned to round $t$; then, for each round $t\in[T]$, it runs an adaptive policy that greedily selects $b_{t}$ items to be observed in round $t$.
It is worth to notice that an optimal solution for SBMSm, in general, dynamically allocates the budget after each item selection. Instead our semi-adaptive approximate policy, at the expense of a small loss in the optimality guarantee, allocates all the budget to the rounds at the beginning.
Remark 1.
Despite our semi-adaptive policy decides the budget allocation in advance, it strongly takes into account the heterogeneity of the input instance over all rounds. Indeed, if the budget was allocated without considering the structure of the input instance (e.g., by distributing uniformly the budget), the resulting approximation could be non-constant.
To show this, assume that the budget allocation is determined by a budget vector $\vec{b}=(b_{1},\ldots,b_{T})$ that depends on the number of rounds $T$ and the budget $B$ only, and let $t^{*}$ be the index $t$ minimizing $b_{t}$ (among $b_{1},b_{2},\ldots,b_{T}$). Now, consider an instance $I=(T,B,V,H,(\mathcal{P}_{t})_{t\in[T]},(f_{t})_{t\in[T]})$ where $|V|=B=T$ and $|H|=1$. As $|H|=1$, the underlying problem is deterministic, thus we can avoid at all the presence of states. Each $f_{t}$ is defined as follows: $f_{t}(S)=|S|$ for any $S\subseteq V$ if $t=t^{*}$ (that is, the state function at round $t^{*}$ counts the number of selected items) and $f_{t}(S)=0$ if $t\neq t^{*}$ (that is, the state function at any round $t\neq t^{*}$ is zeroed).
We have that the optimal policy $\pi^{*}$ for $I$ simply selects all items at round $t^{*}$ (that is, the unique round that produces a positive value of the aggregated state function) and does not select any item at rounds $t\neq t^{*}$; the resulting value of $\pi^{*}$ is $\sigma(\pi^{*})=B=T$. Instead, the value guaranteed by any policy $\pi$ that allocates the budget according to vector $\vec{b}$ is $\sigma(\pi)=b_{t^{*}}=\min_{t\in[T]}b_{t}\leq B/T=T/T=1$. Thus, policy $\pi$ is only $\alpha$-approximate with $\alpha=b_{t^{*}}/B\leq 1/T$ (that is non-constant due to dependence on $T$).
4.1 Estimating State Functions
In order to solve the desired maximization problems, we first need to estimate the expected value of the state functions achieved when selecting an item. However, this is not an easy task due to the stochastic nature of the problem. We first introduce a standard Monte-Carlo algorithm, called ${\sf Oracle1}$, that can be used to estimate the expected value of the state function when selecting an item, with arbitrary precision and arbitrarily high probability: such algorithm can be used to find an item that approximately maximizes the expected influence spread of a state functions, under a given partial state.
Then, we provide another Monte-Carlo algorithm, called ${\sf Oracle2}$, that can be used to estimate the expected increment of the state function achieved by a given adaptive policy $\pi$ when selecting the $j$-th item (in order of selection) at any round $t$, with arbitrary precision and arbitrarily high probability. By exploiting the Chernoff-Hoeffding’s inequality [25], we show that, for any $\delta,\xi>0$ both algorithms can be set to guarantee a precision with an additive error of at most $\delta$ with probability at least $1-\xi$, and to run in polynomial time w.r.t. to the size of the input instance, $1/\delta$ and $1/\xi$. See the SM for a complete description of algorithms ${\sf Oracle1}$, ${\sf Oracle2}$ and their properties.
4.1.1 Oracle1
${\sf Oracle1}$ takes in input an integer $q\geq 1$, an instance $I$ of the SBMSm problem, a round $t\in[T]$, a partial state $\langle S_{t}\rangle$ for round $t$, an item $v$, and picks $q$ samples $X_{t,v,\langle S\rangle}^{1},\ldots,X_{t,v,\langle S\rangle}^{q}$ of the random variable $\bm{X}_{t,v,\langle S\rangle}:=f_{t}(S_{t}\cup\{v\},\bm{\eta}_{t})-f_{t}(S_{t},\bm{\eta}_{t})$ conditioned by the observation of $\langle S_{t}\rangle$. Then, ${\sf Oracle1}$ returns the value ${\sf Oracle1}(q,I,t,\langle S_{t}\rangle,v):=\tilde{\Delta}_{t}(v|\langle S_{t}\rangle):=\sum_{j=1}^{q}X_{t,v,\langle S\rangle}^{j}/q$.
Given $\delta,\xi>0$, ${\sf Oracle1}$ satisfies the following property.
Proposition 2.
If $q\geq q(\delta,\xi):=2\frac{\Lambda^{2}}{\delta^{2}}\ln(2n/\xi)$, then:
  • (i)
    any item $\tilde{v}$ maximizing $\tilde{\Delta}_{t}(\tilde{v}|\langle S_{t}\rangle)$ verifies $\mathbb{P}\left[\max_{v\in V}\Delta_{t}(v|\langle S_{t}\rangle)-{\Delta_{t}}(\tilde{v}|\langle S_{t}\rangle)\leq\delta\right]\geq 1-\xi$;
  • (ii)
    If $q\leq P(q(\delta,\xi))$ for some polynomial $P$, ${\sf Oracle1}$ runs in polynomial time w.r.t. the size of the input instance, $1/\delta$ and $1/\xi$.
Proof.
Let $t\in[T]$, $\langle S\rangle$ be a partial state at round $t$ and, for any $v\in V$, let $\bm{X}_{t,v,\langle S\rangle}$ and $X^{1}_{t,v,\langle S\rangle},\ldots,X^{q}_{t,v,\langle S\rangle}$ be the random quantities considered in the definition of Oracle1. As $\bm{X}_{t,v,\langle S\rangle}$ has values in interval $[0,\Lambda]$ (by the polynomial boundedness of the state function), we have that
$\mathbb{P}\left[\exists v\in V:\Bigg{|}\sum_{j=1}^{q}\frac{X^{j}_{t,v,\langle S\rangle}}{q}-\mathbb{E}\left[\bm{X}_{t,v,\langle S\rangle}\right]\Bigg{|}\geq\delta/2\right]$ $\leq\sum_{v\in V}\mathbb{P}\left[\Bigg{|}\sum_{j=1}^{q}\frac{X^{j}_{t,v,\langle S\rangle}}{q}-\mathbb{E}\left[\bm{X}_{t,v,\langle S\rangle}\right]\Bigg{|}\geq\delta/2\right]$
$\leq n\cdot 2\exp\left(-\frac{2(\delta/2)^{2}q}{\Lambda^{2}}\right)$
$\leq n\cdot 2\exp\left(-\frac{2(\delta/2)^{2}q(\delta,\xi)}{\Lambda^{2}}\right)$
$=\xi.$
where the second inequality follows by the Chernoff-Hoeffding’s inequality, and the last equality holds by construction of $q(\delta,\xi)$. The above inequalities imply that
$\mathbb{P}\left[|\Delta(v|\langle S_{t}\rangle)-\tilde{\Delta}(v|\langle S_{t}\rangle)|\leq\delta/2\ \forall v\in V\right]$ $=\mathbb{P}\left[\Bigg{|}\mathbb{E}\left[\bm{X}_{t,v,\langle S\rangle}\right]-\sum_{j=1}^{q}\frac{X^{j}_{t,v,\langle S\rangle}}{q}\Bigg{|}\leq\delta/2\ \forall v\in V\right]$ (5)
$\geq 1-\xi.$
Now, fix $v^{*}\in\arg\max_{v\in V}\Delta_{t}(v|\langle S_{t}\rangle)$ and $\tilde{v}\in\arg\max_{v\in V}\tilde{\Delta}_{t}(v|\langle S_{t}\rangle)$. For any $q\geq q(\delta,\xi)$, we have that
$\mathbb{P}\left[\max_{v\in V}\Delta(v|\langle S_{t}\rangle)-{\Delta}(\tilde{v}|\langle S_{t}\rangle)\leq\delta\right]$ $=\mathbb{P}\left[\Delta(v^{*}|\langle S_{t}\rangle)-{\Delta}(\tilde{v}|\langle S_{t}\rangle)\leq\delta\right]$
$\geq\mathbb{P}\left[(\Delta(v^{*}|\langle S_{t}\rangle)-\tilde{\Delta}(v^{*}|\langle S_{t}\rangle))+(\tilde{\Delta}(\tilde{v}|\langle S_{t}\rangle)-{\Delta}(\tilde{v}|\langle S_{t}\rangle))\leq\delta\right]$ (6)
$\geq\mathbb{P}\left[|\Delta(v|\langle S_{t}\rangle)-\tilde{\Delta}(v|\langle S_{t}\rangle)|\leq\delta/2\ \forall v\in\{\tilde{v},v^{*}\}\right]$
$\geq\mathbb{P}\left[\Bigg{|}\mathbb{E}\left[\bm{X}_{t,v,\langle S\rangle}\right]-\sum_{j=1}^{q}\frac{X^{j}_{t,v,\langle S\rangle}}{q}\Bigg{|}\leq\delta/2\ \forall v\in V\right]$
$\geq 1-\xi,$ (7)
where (6) holds as $-\tilde{\Delta}(v^{*}|\langle S_{t}\rangle)+\tilde{\Delta}(\tilde{v}|\langle S_{t}\rangle)\geq 0$ (by the optimality of $\tilde{v}$ for $\tilde{\Delta}(\tilde{v}|\langle S_{t}\rangle)$), and (7) follows from (5), and shows property (i).
As for property (ii) we first observe that each oracle sample $X^{j}_{t,v,\langle S_{t}\rangle}$ can be computed in polynomial time. Indeed, for a given partial state $\langle S_{t}\rangle$, we can flip a biased coin with probability $\mu(e)$ for any atom $e\notin R(\langle S_{t}\rangle)$ (i.e., any atom not yet explored by the reachability function $R$), to obtain a state $\eta_{t}\sim\bm{\eta}_{t}$ conditioned by the observation of $\langle S_{t}\rangle$. Then, one can compute $f(S_{t}\cup\{v\},\eta_{t})$ and $f(S_{t},\eta_{t})$ in polynomial time (by exploiting the polynomial time computability of $R$ and $f$) and set $X^{j}_{t,v,\langle S_{t}\rangle}:=f(S_{t}\cup\{v\},\eta_{t})-f(S_{t},\eta)$. Furthermore, $q(\delta,\xi)$ is polynomial, since $\Lambda$ is polynomial in the size of the input. Then $q$ is also polynomial if $q\leq P(q(\delta,\xi))$. ∎
4.1.2 Oracle2
${\sf Oracle2}$ takes in input an integer $q\geq 1$, an instance $I$ of the SBMSm problem, a polynomial adaptive policy $\pi(t)$ that only works at round $t$ and selects all the items in $n$ steps, an integer $i\in[n]$, and picks $q$ samples $Y_{t,i}^{1},\ldots,Y_{t,i}^{q}$ of the random variable $\bm{Y}_{t,i}:=f_{t}(\{\bm{v}_{1},\ldots,\bm{v}_{i}\},\bm{\eta}_{t})-f_{t}(\{\bm{v}_{1},\ldots,\bm{v}_{i-1}\},\bm{\eta}_{t})$, that denotes the increment of $f_{t}$ achieved by $\pi(t)$ when selecting the $i$-th item $\bm{v}_{i}$ at round $t$, where $\bm{v}_{1},\ldots,\bm{v}_{n}$ are the random nodes chosen by $\pi(t)$ in order of selection. Then, ${\sf Oracle2}$ returns the empirical mean ${\sf Oracle2}(q,I,\pi(t),i):=\tilde{\Delta}_{t,i}:=\sum_{j=1}^{q}Y^{j}_{t,i}/q$.
Proposition 3.
If $q\geq q^{\prime}(\delta,\xi):=\frac{\Lambda^{2}}{2\delta^{2}}\ln(2Tn/\xi)$, then:
  • (i)
    $\mathbb{P}\left[|\Delta_{t,i}-\tilde{\Delta}_{t,i}|\leq\delta,\ \forall t\in[T],i\in[n]\right]\geq 1-\xi,$ where $\Delta_{t,i}$ is the (exact) expected increment of $f_{t}$ achieved by $\pi(t)$ when selecting the $i$-th item;
  • (ii)
    if $q\leq P(q^{\prime}(\delta,\xi))$ for some polynomial $P$, ${\sf Oracle2}$ runs in polynomial time w.r.t. the size of the input instance, $1/\delta$ and $1/\xi$.
Proof.
Given $t\in[T]$, $i\in[n]$ and $q\geq 1$, let $\bm{Y}_{t,i}$ and $Y_{t,i}^{1},\ldots,Y_{t,i}^{q}$ be the random quantities considered in the definition of Oracle2. As $\bm{Y}_{t,i}$ has values in $[0,\Lambda]$ we have that:
$\mathbb{P}\left[\exists t\in[T],i\in[n]:\Bigg{|}\sum_{j=1}^{q}\frac{Y^{j}_{t,i}}{q}-\mathbb{E}\left[\bm{Y}_{t,i}\right]\Bigg{|}\geq\delta\right]$ $\leq\sum_{t\in[T]}\sum_{i\in[n]}\mathbb{P}\left[\Bigg{|}\sum_{j=1}^{q}\frac{X^{j}_{t,v,\langle S\rangle}}{q}-\mathbb{E}\left[\bm{X}_{t,v,\langle S\rangle}\right]\Bigg{|}\geq\delta\right]$
$\leq T\cdot n\cdot 2\exp\left(-\frac{2\delta^{2}q}{\Lambda^{2}}\right)$
$\leq T\cdot n\cdot 2\exp\left(-\frac{2\delta^{2}q^{\prime}(\delta,\xi)}{\Lambda^{2}}\right)$
$=\xi.$
where the second inequality holds by the Chernoff-Hoeffding’s inequality, and the last equality holds by construction of $q^{\prime}(\delta,\xi)$. The above inequalities imply that
$\mathbb{P}\left[|\tilde{\Delta}^{\prime}_{t,i}-\Delta^{\prime}_{t,i}|\leq\delta,\ \forall t\in[T],i\in[n]\right]$ $=\mathbb{P}\left[\Bigg{|}\sum_{j=1}^{q}\frac{Y^{j}_{t,i}}{q}-\mathbb{E}\left[\bm{Y}_{t,i}\right]\Bigg{|}\leq\delta,\ \forall t\in[T],i\in[n]\right]$
$\geq 1-\xi,$
and this shows property (i).
Property (ii) holds by similar arguments as in Proposition 2. ∎
4.2 Single-Round Adaptive Policy
In order to define our approximate policy we first introduce the greedy framework adopted at each round, given the available budget for that round. The greedy single-round (adaptive) policy ${\sf SingleGr}_{\delta,\xi}(I,t,b)$ (see Algorithm 1 for the pseudo-code) greedily selects $b$ items (in $b$ steps) at round $t$ only, and at each step selects an item maximizing the expected increment of $f$ up to an addend $\delta$ and with prob. at least $1-\xi$, conditioned by the previous observations; the item selected at each step can be determined by calling Oracle1 for each item $v$ not yet included, and then choosing that maximizing the value ${\sf Oracle1}(q(\delta,\xi),I,t,\langle S_{t}\rangle,v)$.
Algorithm 1 ${\sf SingleGr}_{\delta,\xi}$
0:  an input instance $I$ of SBMSm, a round $t$, a budget $b\geq 0$, an (unknown) state $\eta_{t}$ for round $t$;
0:  a subset $S_{t}:={\sf SingleGr}_{\delta,\xi}(I,t,b,\eta_{t})$;
1:  let $S_{t}\leftarrow\emptyset$;
2:  while $|S_{t}|\leq b$ do
3:     $v\leftarrow\arg\max_{v^{\prime}\in V}{\sf Oracle1}(q(\delta,\xi),I,t,\langle S_{t}\rangle,v^{\prime})$;$\Delta_{t}(v|\langle S_{t},\eta_{t}\rangle)\geq\max_{v^{\prime}\in V\setminus S_{t}}\Delta_{t}(v^{\prime}|\langle S_{t},\eta_{t}\rangle)-\delta$ with prob. at least $1-\xi$ by Proposition 2 (see Section 4.1.1)
4:     $S_{t}\leftarrow S_{t}\cup\{v\}$;
5:  end while
6:  return  $S_{t}$.
Notice that ${\sf SingleGr}_{\delta,\xi}$ is equivalent to an approximate variant of the adaptive greedy policy defined by [22], that provides a $\left(1-\frac{1}{e}\right)$-approximation to the best policy when considering a unique round.
4.3 Budget Allocation Algorithm
In order to allocate the budget over rounds, we consider an algorithm called ${\sf BudgetGr}_{\delta,\epsilon}(I)$ (see Algorithm 2 for the pseudo-code), that greedily assigns (in a non-adaptive way) each unit of budget to the round $t$ that maximizes the expected increment of the state function achieved by the greedy single-round (adaptive) policy; the above expected increment is measured with arbitrarily precision and arbitrarily high probability by algorithm Oracle2, in which the single-greedy policy is executed many times to collect a sufficiently high number of samples on the state function increments.
Algorithm 2 ${\sf BudgetGr}_{\delta,\xi}$
0:  an input instance $I$ of SBMSm;
0:  a vector $\vec{b}:=(b_{1},\ldots,b_{T})={\sf BudgetGr}_{\delta,\xi}(I)$;
1:  for $t=1,\ldots,T$ do
2:     $\overline{\Delta}_{t,0}\leftarrow\infty$;
3:     for $i=1,\ldots,n$ do
4:        $\tilde{\Delta}_{t,i}\leftarrow{\sf Oracle2}(q^{\prime}(\delta,\xi),I,{\sf SingleGr}_{\delta,\xi}(I,t,n),i)$;
5:        $\overline{\Delta}_{t,i}\leftarrow\min\{\overline{\Delta}_{t,i-1},\tilde{\Delta}_{t,i}\}$;
6:     end for
7:     $\overline{\Delta}_{t,n+1}\leftarrow-\infty$;
8:  end for
9:  % By Proposition 3 (see Section 4.1.2) applied to ${\sf SingleGr}_{\delta,\xi}(I,t,n)$, “$|\Delta_{t,i}-\tilde{\Delta}_{t,i}|\leq\delta,\ \forall t\in[T],i\in[n]$” holds with prob. at least $1-\xi$, where $\Delta_{t,i}$ denotes the (exact) expected increment of state function $f_{t}$, achieved by ${\sf SingleGr}_{\delta,\xi}(I,t,n)$ when selecting the $i$-th item, in order of selection. Furthermore, as $\Delta_{t,i}$ is non-increasing in $i$ up to a small addend (by adaptive-submodularity), we also have that the distance $|\Delta_{t,i}-\overline{\Delta}_{t,i}|$ is small with prob. at least $1-\xi$.
10:  $(b_{1},\ldots,b_{T})\leftarrow(0,\ldots,0)$;
11:  for $h=1,\ldots,B$ do
12:     $t^{*}\leftarrow\arg\max_{t\in[T]}\overline{\Delta}_{t,b_{t}+1}$;
13:     $b_{t^{*}}\leftarrow b_{t^{*}}+1$;
14:  end for
15:  return  $\vec{b}:=(b_{1},\ldots,b_{T})$;
4.4 Multi-Round Adaptive Policy
Let $\vec{b}:=(b_{1},\ldots,b_{T})$ be the output of ${\sf BudgetGr}_{\delta,\epsilon}(I)$. The greedy multi-round (adaptive) policy ${\sf MultiGr}_{\delta,\xi}(I)$ (see Algorithm 1) applies the greedy single-round policy with budget $b_{t}$ at each round $t\in[T]$ to determine a subset $S_{t}$, and returns the aggregated subset $\vec{S}=(S_{1},\ldots,S_{T})$.
Algorithm 3 ${\sf MultiGr}_{\delta,\xi}$
0:  an input instance $I$ of SBMSm, an (unknown) aggregated state $\vec{\eta}:=(\eta_{1},\ldots,\eta_{t})$;
0:  an aggregated subset $\vec{S}:={\sf MultiGr}_{\delta,\xi}(I,\vec{\eta})$;
1:  $\vec{b}:=(b_{1},\ldots,b_{T})\leftarrow{\sf BudgetGr}_{\delta.\xi}(I)$;
2:  for $t=1,\ldots,T$ do
3:     $S_{t}\leftarrow{\sf SingleGr}_{\delta,\xi}(I,t,b_{t},\eta_{t})$;
4:     wait the end of round $t$;
5:  end for
6:  return  $\vec{S}:=(S_{1},\ldots,S_{T})$.
Theorem 1 shows that ${\sf MultiGr}_{\delta,\xi}(I)$ is a $(1/2,\epsilon)$-approximate policy for any $\epsilon>0$, for opportune choices of $\delta$ and $\xi$.
Theorem 1.
Let $I=(T,B,V,H,(\mathcal{P}_{t})_{t\in[T]},R,(f_{t})_{t\in[T]})$ be an input instance of SBMSm. The greedy multi-round policy ${\sf MultiGr}_{\delta,\xi}(I)$ is a $\left(1/2-\epsilon\right)$-approximate policy for $\delta:=\xi:=\frac{\lambda c\epsilon}{B(1+\Lambda)}$ and any constant $c$ such that $40<c\leq 1/$ 4, and it is polynomial in the size of $I$ and $1/\epsilon$.
Before to provide the full proof of this theorem (provided in the next section), we would like to describe the main proof arguments necessary to get a $1/2$-approximation assuming that the probabilistic oracles are exact (i.e., assuming that Oracle1 and Oracle2 can be run with $\delta:=\xi:=0$).
Consider the following artificial adaptive policy $\hat{\pi}^{*}$ that works in two phases: in the first phase $\hat{\pi}^{*}$ selects $B$ items by running the greedy multi-round policy ${\sf MultiGr}_{0,0}(I)$; in the second phase, $\hat{\pi}^{*}$ restarts from round $1$, and by observing the partial state of the previous item selections, it selects further $\min\{nT-B,B\}$ items in an optimal adaptive way over $T$ rounds. By construction, the performance of $\hat{\pi}^{*}$ is not worse than that of an optimal adaptive policy $\pi^{*}$ selecting $B$ items only. By exploiting the adaptive submodularity, one can show that the expected increment of the state functions achieved by any item selected in the first phase is at least equal as that achieved by each item selected in the second phase. Thus, considering that at most $B$ items are chosen in the second phase, and exactly $B$ items are chosen in the first one, we necessarily have that the expected increment of the aggregated state function achieved in the first phase (that is, $\sigma({\sf MultiGr}_{0,0}(I))$) is at least equal as that achieved in the second phase (that is, $\sigma(\hat{\pi}^{*})-\sigma({\sf MultiGr}_{0,0}(I))$). Then, by rearranging the previous inequality, we get $\sigma({\sf MultiGr}_{0,0}(I))\geq\frac{1}{2}\sigma(\hat{\pi}^{*})\geq\frac{1}{2}\sigma(\hat{\pi})$, and this shows the desired approximation.
4.5 Full Proof of Theorem 1
First of all, the polynomial time complexity follows from the fact that the greedy multi-round policy applies $n\cdot T$ times algorithm Oracle2 to compute the budget vector (via algorithm ${\sf BudgetGr}_{\delta,\xi}(I)$) and $B<n\cdot T$ times the algorithm Oracle1 to select $B$ items, and since both Oracle1 and Oracle2 run in polynomial time. Thus, in the remainder of the proof we will only focus on the approximation guarantee of the greedy multi-round policy.
Fix $\delta:=\xi:=\frac{\lambda c\epsilon}{B(1+\Lambda)}$, where $c\in(0,1/4]$ is a constant that can be arbitrarily fixed. Let $\pi^{*}$ denote an optimal adaptive policy for instance $I$, and consider the following artificial adaptive policy $\hat{\pi}^{*}$ that works in two phases. Let ${\sf BudgetGr}^{\prime}_{\delta,\xi}(I)$ be a variant of ${\sf BudgetGr}_{\delta,\xi}(I)$ working assuming that $\tilde{\Delta}_{t,i}=\Delta_{t,i}$ holds with probability $1$ (that is, such variant computes the budget vector based on the knowledge of the exact expected increments of the state functions), and let $\vec{b}^{\prime}=(b_{1}^{\prime},\ldots,b_{T}^{\prime})$ be the resulting budget vector. In the first phase, $\hat{\pi}^{*}$ executes a variant ${\sf MultiGr}^{\prime}_{\delta,\xi}(I)$ of the greedy multi-round policy that uses the budget vector $\vec{b}^{\prime}=(b_{1}^{\prime},\ldots,b_{T}^{\prime})$ to distribute the budget among all $T$ rounds, and let $\vec{S}=(S_{1},\ldots,S_{T})$ denote the aggregated set (with $\sum_{t\in[T]}|S_{t}|=B$) returned by ${\sf MultiGr}^{\prime}_{\delta,\xi}(I)$. In the second phase, $\pi^{*}$ restarts from round $1$, and by observing the aggregated partial state $\langle\vec{S}\rangle$ it selects further $\min\{nT-B,B\}$ items in an optimal adaptive way over $T$ rounds.
By construction of $\hat{\pi}^{*}$, we can easily observe that the expected performance of policy $\hat{\pi}^{*}$ is not worse than that of $\pi^{*}$, that is
$\sigma(\hat{\pi}^{*})\geq\sigma(\pi^{*}).$ (8)
In the following lemma, by exploiting the adaptive submodularity, we show that the expected value of the aggregated state function achieved in the first phase (that is, $\sigma({\sf MultiGr}^{\prime}_{\delta,\xi}(I))$) is at least equal to the expected increment of the aggregated state function achieved in the second phase (that is, $\sigma(\hat{\pi}^{*})-\sigma({\sf MultiGr}^{\prime}_{\delta,\xi}(I))$), up to an addend $B(\delta+\xi\Lambda)$.
Lemma 1.
We have $\sigma({\sf MultiGr}^{\prime}_{\delta,\xi}(I))\geq\sigma(\hat{\pi}^{*})-\sigma({\sf MultiGr}^{\prime}_{\delta,\xi}(I))-2B(\delta+\Lambda\xi)$.
Proof.
Given a round $t\in[T]$ and $j\in[b_{t}^{\prime}]$, let $\bm{v}_{t,i}$ denote the (random) $i$-th item selected at round $t$ by the greedy multi-round policy ${\sf MultiGr}_{\delta,\xi}(I)$ (when it applies, as sub-routine, the greedy single-round policy) and let $\bm{S}_{t,i}$ denote the (random) set $\{\bm{v}_{t,1},\ldots,\bm{v}_{t,i}\}$. We recall that $\Delta_{t,i}=\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}[\Delta(\bm{v}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)]$ denotes the expected increment of the state function at round $t$ achieved by ${\sf SingleGr}_{\delta,\xi}(I,t,n)$ when selecting the $i$-th item at round $t$; furthermore, let $\bm{v}_{t,i}^{*}$ be the (random) item that maximizes $\Delta(\bm{v}^{*}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)$ and let $\Delta^{*}_{t,i}:=\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}[\Delta(\bm{v}^{*}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)]$.
In the remainder of the proof, we will first show several inequalities (Parts i-iv) and then we will combine them to show the claim (Final Part).
Part i:
Given a round $t\in[T]$ and $i\in[b_{t}^{\prime}]$, let $A_{t,i}$ denote the event “$\Delta(\bm{v}^{*}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)\leq\Delta(\bm{v}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)+\delta$”. By the property (i) of Proposition 2 (see Section 4.1.1), we have that event $A_{t,i}$ is true with probability at least $1-\xi$, and then, its complementary event $\overline{A_{t,i}}$ is false with probability at most $\xi$. Thus, we have
$\Delta^{*}_{t,i}$ $=\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}[\Delta(\bm{v}^{*}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)]$ (9)
$=\mathbb{P}[A_{t,i}]\cdot\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}[\Delta(\bm{v}^{*}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)\ |\ A_{t,i}]+\mathbb{P}\left[\overline{A_{t,i}}\right]\cdot\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}\left[\Delta(\bm{v}^{*}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)\ |\ \overline{A_{t,i}}\right]$
$\leq\mathbb{P}[A_{t,i}]\cdot\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}[\Delta(\bm{v}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)+\delta\ |\ A_{t,i}]+\mathbb{P}\left[\overline{A_{t,i}}\right]\cdot\Lambda$
$\leq\mathbb{P}[A_{t,i}]\cdot\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}[\Delta(\bm{v}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)+\delta\ |\ A_{t,i}]+\xi\Lambda$
$\leq\mathbb{P}[A_{t,i}]\cdot\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}[\Delta(\bm{v}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)+\delta\ |\ A_{t,i}]+\mathbb{P}[\overline{A}_{t,i}]\cdot\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}[\Delta(\bm{v}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)+\delta\ |\ \overline{A}_{t,i}]+\xi\Lambda$
$=\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}[\Delta(\bm{v}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)+\delta]+\xi\Lambda$
$=\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle}[\Delta(\bm{v}_{t,i}|\langle\bm{S}_{t,i-1}\rangle)]+\delta+\xi\Lambda$
$=\Delta_{t,i}+\delta+\xi\Lambda.$
Furthermore, given a round $t\in[T]$ and $i,i^{\prime}\in[b_{t}^{\prime}]$ with $i^{\prime}\leq i$, by exploiting the adaptive submodularity we have that:
$\Delta_{t,i^{\prime}}^{*}$ $=\mathbb{E}_{\langle\bm{S}_{t,i^{\prime}-1}\rangle}[\Delta(\bm{v}_{t,i^{\prime}}^{*}|\langle\bm{S}_{t,i^{\prime}-1}\rangle)]$ (10)
$\geq\mathbb{E}_{\langle\bm{S}_{t,i^{\prime}-1}\rangle,\bm{v}_{t,i}^{*}}[\Delta(\bm{v}_{t,i}^{*}|\langle\bm{S}_{t,i^{\prime}-1}\rangle)]$ (11)
$\geq\mathbb{E}_{\langle\bm{S}_{t,i-1}\rangle,\bm{v}^{*}_{t,i}}^{*}[\Delta(\bm{v}_{t,i}^{*}|\langle\bm{S}_{t,i-1}\rangle)]$ (12)
$=\Delta^{*}_{t,i},$ (13)
where (11) holds by the optimality of $\bm{v}_{t,i}^{*}$ and (12) holds by adaptive submodularity.
Part ii:
Now, for any $j\in[B]$, let $(t_{j},i_{j})$ be the pair such that $t_{j}$ is the round to which the $j$-th unit of budget is assigned by ${\sf BudgetGr}^{\prime}_{\delta,\xi}(I)$, and $i_{j}$ denotes the budget used at round $t_{j}$ after assigning the $j$-th unit of budget. Given $j\in[B]$, we have that
$\Delta_{t_{B},i_{B}}$ $=\Delta_{t_{B},b^{\prime}_{t_{B}}}$ (14)
$\leq\Delta_{t_{j},b^{\prime}_{t_{j}}}$ (15)
$\leq\Delta^{*}_{t_{j},b^{\prime}_{t_{j}}}$
$\leq\Delta^{*}_{t_{j},i_{j}}$ (16)
$\leq\Delta_{t_{j},i_{j}}+\delta+\xi\Lambda,$ (17)
where (14) holds since $i_{B}$ is equal to the budget assigned at round $t$ at the end of algorithm ${\sf BudgetGR}^{\prime}_{\delta,\xi}(I)$, (15) holds by the greedy choice of ${\sf BudgetGR}^{\prime}_{\delta,\xi}(I)$, (16) follows from (13) and (17) follows from (9).
By exploiting (17) we get
$\sigma({\sf MultiGr}^{\prime}_{\delta,\xi}(I))$ $=\sum_{j\in[B]}\Delta_{t_{j},i_{j}}$ (18)
$\geq\sum_{j\in[B]}\left(\Delta^{*}_{t_{j},i_{j}}-(\delta+\xi\Lambda)\right)$
$\geq\sum_{j\in[B]}\left(\Delta^{*}_{t_{B},i_{B}}-(\delta+\xi\Lambda)\right)$
$=B\cdot\Delta^{*}_{t_{B},i_{B}}-B(\delta+\xi\Lambda),$
where the first inequality follows from (9) and the second one from (13).
Part iii:
Finally, for any $h\in[\min\{B,Tn-B\}]$, let $\hat{\Delta}^{*}_{h}$ denote the expected increment of the $h$-th item selected by $\hat{\pi}^{*}$ since the beginning of the second phase (in this case, the ordering is not related to the specific round). By construction, we have that $\sum_{h\in[\min\{B,Tn-B\}]}\hat{\Delta}^{*}_{h}$ is equal to the expected increment achieved in the second phase of $\hat{\pi}^{*}$, that is,
$\sum_{h\in[\min\{B,Tn-B\}]}\hat{\Delta}^{*}_{h}=\sigma(\hat{\pi}^{*})-\sigma({\sf MultiGr}^{\prime}_{\delta}(I)).$ (19)
For a given $h\in[\min\{B,Tn-B\}]$, let $t$ be the round in which the $h$-th item is assigned. We know that $b_{t}$ items have been previously selected in the first phase, when the greedy multi-round policy was run. Thus, by exploiting again the adaptive submodularity, we have that the expected increment $\Delta^{*}_{t,b_{t}+1}$ (i.e., that obtained by selecting optimally an item in round $t$, at the end of the first phase) is at least equal as that achieved by $\hat{\pi}^{*}$ by adding the $h$-th item in the second phase, that is:
$\hat{\Delta}^{*}_{h}\leq\Delta^{*}_{t,b_{t}+1}\leq\Delta_{t,b_{t}+1}+\delta+\xi\Lambda\leq\Delta_{t_{B},i_{B}}+\delta+\xi\Lambda,$ (20)
where the first inequality holds by the above observations, the second one holds by (13) and the last one holds by the greedy choice of ${\sf BudgetGR}^{\prime}_{\delta,\xi}(I)$.
Final Part:
By putting together (18), (19) and (20), we obtain
[…]
$\geq B\cdot\Delta_{t_{B},i_{B}}-B(\delta+\xi\Lambda)$
$\geq\min\{B,Tn-B\}\cdot\Delta_{t_{B},i_{B}}-B(\delta+\xi\Lambda)$
[…]
$\geq\sum_{h\in[\min\{B,Tn-B\}]}\hat{\Delta}^{*}_{h}-2B(\delta+\xi\Lambda)$
[…]
and this shows the claim of the lemma. ∎ In the following lemma, we relate the value of ${\sf MultiGr}_{\delta,\xi}(I)$ to that of its variant ${\sf MultiGr}^{\prime}_{\delta,\xi}(I)$.
Lemma 2.
We have $\sigma({\sf MultiGr}_{\delta,\xi}(I))\geq\sigma({\sf MultiGr}^{\prime}_{\delta,\xi}(I))-3B(\delta+\xi\Lambda)$.
Proof.
As in Lemma 1, we will first show several inequalities (Parts i-v), and then we will combine them to show the claim (Final Part).
Part i:
We recall that $\Delta_{t,i}$ denotes the expected increment of the state function considered in algorithm ${\sf SingleGr}_{\delta,\xi}(I,t,n)$, when selecting the $i$-the item at round $t$; by construction of policy ${\sf MultiGr}^{\prime}_{\delta,\xi}(I)$, we have that
$\sigma({\sf MultiGr}^{\prime}_{\delta,\xi}(I))=\sum_{t\in[T]}\sum_{i\in[b_{t}^{\prime}]}\Delta_{t,i}.$ (21)
Let $\vec{\bm{b}}=(\bm{b}_{1},\ldots,\bm{b}_{T})$ be the (random) vector returned by ${\sf BudgetGr}_{\delta,\xi}(I)$; by construction, we have that
$\sigma({\sf MultiGr}_{\delta,\xi}(I))=\mathbb{E}\left[\sum_{t\in[T]}\sum_{i\in[\bm{b}_{t}]}{\Delta}_{t,i}\right].$ (22)
Part ii:
By exploiting similar arguments as in inequality (17) of Lemma 1, one can show that the $\Delta_{t,i}$ is non-increasing in $i$ up to addend $(\delta+\xi\Lambda)$, that is,
$\Delta_{t,i^{\prime}}\geq\Delta_{t,i}-(\delta+\xi\Lambda)$ (23)
for any $t\in[T]$ and $i,i^{\prime}\in[n]$ such that $i^{\prime}\leq i$; in particular, such inequality can be shown by observing that $\Delta_{t,i^{\prime}}+\delta+\xi\Lambda\geq\Delta_{t,i^{\prime}}^{*}\geq\Delta_{t,i}^{*}\geq\Delta_{t,i}$, where $\Delta_{t,j}^{*}$ is the maximum expected increment when adding an item at round $t$, assuming that $j-1$ items have been already selected in that round by the approximate greedy single-round policy (see the proof of Lemma 1 for a formal definition of $\Delta_{t,j}^{*}$).
Part iii:
Now, let $\tilde{\bm{\Delta}}_{t,i}$ be the random quantity considered in algorithm Oracle2 to estimate $\Delta_{t,i}$ with high probability and arbitrarily high precision. Let $A$ be the event “$|\Delta_{t,i}-\tilde{\bm{\Delta}}_{t,i}|\leq\delta,\ \forall t\in[T],i\in[n]$”; by the property (i) of Proposition 3 (see Section 4.1.1), we have that $\mathbb{P}[A]\geq 1-\xi$ and $\mathbb{P}[\overline{A}]=1-\mathbb{P}[A]\leq\xi$, where $\overline{A}$ denotes the complementary event of $A$. Then, given a realization $(\tilde{\Delta}_{t,i})_{t\in[T],i\in[n]}\sim(\tilde{\bm{\Delta}}_{t,i})_{t\in[T],i\in[n]}$ such that event $A$ is true we have that
$|\Delta_{t,i}-\tilde{\Delta}_{t,i}|\leq\delta.$ (24)
Part iv:
Let $\overline{\bm{\Delta}}_{t,i}=\min_{i^{\prime}\in[i]}\tilde{\bm{\Delta}}_{t,i^{\prime}}$ be the (random) values considered in ${\sf BudgetGr}_{\delta,\xi}(I)$ to assign the budget in a greedy way. Given a realization $(\overline{\Delta}_{t,i})_{t\in[T],i\in[n]}\sim(\overline{\bm{\Delta}}_{t,i})_{t\in[T],i\in[n]}$ of the above values, and the corresponding budget vector $\vec{b}=(b_{1},\ldots,b_{T})$ returned as output by ${\sf BudgetGr}_{\delta,\xi}(I)$, we have that
$\sum_{t\in[T]}\sum_{i\in[b_{t}]}\overline{\Delta}_{t,i}\geq\sum_{t\in[T]}\sum_{i\in[b_{t}^{\prime}]}\overline{\Delta}_{t,i}.$ (25)
Indeed, as $\overline{\Delta}_{t,i}$ is non increasing in $i$ (by definition of minimum), one can easily show that the vector $\vec{b}^{*}=(b_{1}^{*},\ldots,b_{T}^{*})$ maximizing $\sum_{i\in[b_{t}^{*}]}\overline{\Delta}_{t,i}$ subject to $\sum_{t\in[T]}b_{t}^{*}=B$ can be computed in a greedy way as in algorithm ${\sf BudgetGr}_{\delta,\xi}(I)$, that is, the budget vector $\vec{b}$ returned by ${\sf BudgetGr}_{\delta,\xi}(I)$ is optimal and then (25) holds.
Part v:
Let $(\tilde{\Delta}_{t,i})_{t\in[T],i\in[n]}$ be a realization of $(\tilde{\bm{\Delta}}_{t,i})_{t\in[T],i\in[n]}$ such that event $A$ is true, and let $\overline{\Delta}_{t,j}:=\min_{i^{\prime}\in[i]}\tilde{\Delta}_{t,i}$. Given $t\in[T],i\in[n]$, let $i^{\prime}\in[i]$ be the lowest integer such that $\tilde{\Delta}_{t,i^{\prime}}=\overline{\Delta}_{t,i}$. By exploiting (23), we get:
$\Delta_{t,i}-\overline{\Delta}_{t,i}$ $=\Delta_{t,i}-\tilde{\Delta}_{t,i^{\prime}}$
$\leq\Delta_{t,i^{\prime}}-\tilde{\Delta}_{t,i^{\prime}}+\delta+\xi\Lambda$ (26)
$\leq 2\delta+\xi\Lambda,$ (27)
where (26) holds by (23) and (27) holds by (24).
Final Part:
We conclude the proof by using all the above inequalities as follows:
$\sigma({\sf MultiGr}^{\prime}_{\delta,\xi}(I))-\sigma({\sf MultiGr}_{\delta,\xi}(I))$ (28)
$=\sum_{t\in[T]}\sum_{i\in[b_{t}^{\prime}]}\Delta_{t,i}-\mathbb{E}\left[\sum_{t\in[T]}\sum_{i\in[\bm{b}_{t}]}{\Delta}_{t,i}\right]$
$=\sum_{t\in[T]}\sum_{i\in[b_{t}^{\prime}]}\Delta_{t,i}-\mathbb{E}\left[\sum_{t\in[T]}\sum_{i\in[b_{t}^{\prime}]}\overline{\bm{\Delta}}_{t,i}\right]$
$\ \ \ \ \ \ \ +\mathbb{E}\left[\sum_{t\in[T]}\sum_{i\in[b_{t}^{\prime}]}\overline{\bm{\Delta}}_{t,i}\right]-\mathbb{E}\left[\sum_{t\in[T]}\sum_{i\in[\bm{b}_{t}]}\overline{\bm{\Delta}}_{t,i}\right]+\mathbb{E}\left[\sum_{t\in[T]}\sum_{i\in[\bm{b}_{t}]}\overline{\bm{\Delta}}_{t,i}\right]-\mathbb{E}\left[\sum_{t\in[T]}\sum_{i\in[\bm{b}_{t}]}\tilde{\bm{\Delta}}_{t,i}\right]$
$\ \ \ \ \ \ \ +\mathbb{E}\left[\sum_{t\in[T]}\sum_{i\in[\bm{b}_{t}]}\tilde{\bm{\Delta}}_{t,i}\right]-\mathbb{E}\left[\sum_{t\in[T]}\sum_{i\in[\bm{b}_{t}]}{\Delta}_{t,i}\right]$
$=\mathbb{E}\left[\sum_{t\in[T]}\sum_{i\in[b_{t}^{\prime}]}(\Delta_{t,i}-\overline{\bm{\Delta}}_{t,i})\right]$
[…]
$\ \ \ \ \ \ \ +\mathbb{E}\left[\sum_{t\in[T]}\sum_{i\in[\bm{b}_{t}]}(\tilde{\bm{\Delta}}_{t,i}-\Delta_{t,i})\right]$
$\leq\mathbb{E}\left[\sum_{t\in[T]}\sum_{i\in[b_{t}^{\prime}]}(\Delta_{t,i}-\overline{\bm{\Delta}}_{t,i})\right]+\mathbb{E}\left[\sum_{t\in[T]}\sum_{i\in[\bm{b}_{t}]}(\tilde{\bm{\Delta}}_{t,i}-\Delta_{t,i})\right]$
[…]
$\leq B(2\delta+\xi\Lambda)+B\xi\Lambda+B\delta+B\xi\Lambda$
$=3B(\delta+\xi\Lambda).$
By (28), the claim follows. ∎
Lemma 3.
We have $\lambda\leq\sigma(\pi^{*})$.
Proof.
By definition of $\lambda$, there exists $t\in[T]$ and $v\in V$ such that $\lambda\leq\mathbb{E}_{\bm{\eta}_{t}\sim\mathcal{P}_{t}}[f_{t}(\{v\},\bm{\eta}_{t})]$. By the optimality of policy $\pi^{*}$, we necessarily have $\mathbb{E}_{\bm{\eta}_{t}\sim\mathcal{P}_{t}}[f_{t}(\{v\},\bm{\eta}_{t})]\leq\sigma(\pi^{*})$. Thus, by combining the obtained inequalities, the claim of the lemma follows. ∎
By combining Lemma 1 with (8), we get $\sigma({\sf MultiGr}^{\prime}_{\delta,\xi}(I))\geq\sigma(\pi^{*})-\sigma({\sf MultiGr}^{\prime}_{\delta,\xi}(I))-2B(\delta+\Lambda\xi)$, that is
$\sigma({\sf MultiGr}^{\prime}_{\delta,\xi}(I))\geq\frac{\sigma(\pi^{*})}{2}-B(\delta+\Lambda\xi).$ (29)
Then, we get
$\sigma({\sf MultiGr}_{\delta,\xi}(I))$ $\geq\sigma({\sf MultiGr}^{\prime}_{\delta,\xi}(I))-3B(\delta+\xi\Lambda)$ (30)
$\geq\frac{1}{2}\sigma(\pi^{*})-B(\delta+\Lambda\xi)-3B(\delta+\xi\Lambda)$ (31)
$=\frac{1}{2}\sigma(\pi^{*})-4\delta(1+\Lambda)B$
$\geq\frac{1}{2}\sigma(\pi^{*})-\frac{\delta(1+\Lambda)B}{c}$
$\geq\frac{1}{2}\sigma(\pi^{*})-\epsilon\cdot\lambda$ (32)
$\geq\frac{1}{2}\sigma(\pi^{*})-\epsilon\cdot\sigma(\pi^{*})$ (33)
$=\left(\frac{1}{2}-\epsilon\right)\sigma(\pi^{*}),$ (34)
where (30) follows from Lemma 2, (31) follows from (29), (32) holds since we previously set $\delta:=\xi:=\frac{\lambda c\epsilon}{B(1+\Lambda)}$ with $c\leq 1/4$, and (33) holds by Lemma 3. By (34), the theorem follows.
5 Conclusions
In this paper we presented the first results for the problem of adaptively optimizing the value of a stochastic submodular function over multiple rounds with total budget constraints when the state distributions are known.
Our analysis aims to provide theoretical guarantees on the approximation that can be achieved in polynomial time. We proposed two algorithms: one based on a dynamic programming approach, that is able to compute the optimal multi-round adaptive policy in polynomial time whenever the single-round optimization problem can be efficiently solved, and a simpler greedy algorithm that, despite being only semi-adaptive, still returns a constant approximation of the optimum. Even if the proposed algorithms are essentially simple (mainly the second one that follows a greedy programming technique) and they are proved to run in polynomial time, they involve some routines, such as Monte Carlo simulations, that are known to limit their practical application. Anyway, we believe that, as occurred for many similar problems, our algorithms may foster further research towards more practical approaches, that, e.g., replace the more expensive routines with heuristics. Not only, but our algorithm will also serve as a benchmark for the evaluation of these heuristics.
Apart for optimizing the algorithm for practical application, a theoretically relevant open question is that of providing better approximation algorithms for the multi-round optimization problem studied in this work. We point out that, for submodular single-round optimization problems under budget constraints, there is a $(1-1/e-\epsilon)$ approximation algorithm [22], and cannot be substantially improved in light of the $(1-1/e+\epsilon)$-inapproximability shown by [19]. Thus, it would be interesting to close the gap between $(1-1/e-\epsilon)$ and $(1-1/2-\epsilon)$ on the approximation to the multi-round optimization problem.
Finally, we recall that our algorithms assume that state distributions are known in advance, even if their realization is unknown. It would be undoubtedly interesting to explore the existence of algorithms with good approximation guarantees even if the state distributions must be learnt or they are provided by a some learning framework (without any precision guarantee).